Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

Sales $ 40,000
Variable expenses 26,000
Contribution margin 14,000
Fixed expenses 8,680
Net operating income $ 5,320

9. What is the break-even point in dollar sales?

10. How many units must be sold to achieve a target profit of $8,400?

11. What is the margin of safety in dollars? What is the margin of safety percentage?

margin of safety in dollars

margin of safety percentage

12. What is the degree of operating leverage? (Round your answer to 2 decimal places.)
Answer #1

Break even sales = Fixed cost/Contribution margin ratio

Contribution margin ratio = 14000/40000 = 35%

= 8680/35%

= 24,800

Units to be sold = (Fixed costs+target profit) /Contribution margin per unit

= (8680+8400)/14

= 1220

Margin of Safety = Sales - Breakeven sales

= 40,000 - 24,800

= 15,200

Margin of Safety percentage = 15,200/40,000

= 38%

Degree of operating leverage = Contribution margin/net operating income

= 14,000/5320

= 2.63